OpenStudy (anonymous):

find the quadratic equation from, 1/2xsqrt1-x^2=1.2, has roots 0.88 and 0.469

OpenStudy (tkhunny):

You must make a little more effort to write clearly. Use parentheses to clarify. Is it \[\frac{1}{2}x\cdot\sqrt{1-x^{2}} = 1.2\]

OpenStudy (anonymous):

sorry, my first time still learning how it all works, how do i get from 1/2x*sqrt1-x^2 into ax+bx+c where i can solve using quadratic formula, really appreciate your help,

OpenStudy (tkhunny):

First, no worries. Just continue to take encouragement to learn to communicate better. This faceless communication can be a little tricky. Second, you don't. You must first remember where it came from. In this case, that square root it a little bit of trouble. 1 - x^2 must not be negative. Do you know why?

OpenStudy (anonymous):

Not sure, it was an equation that came via an civil engineering tutorial that i was doing in relation to vibration analysis ( dynamic magnification ), i'll try to give the original equation, hopefully do a better job this time. 1/2zi*sqrt(1-zi^2) it's a tricky one, if you can't help that ok, appreciate what you have done, regards PQ.

OpenStudy (tkhunny):

It's okay. Generally, we want to stay in Real Numbers. Square roots of negative numbers are a little tricky if you want to stary in Real Numbers. Rewrite a little: \[\sqrt{1-x^{2}} = \frac{1.2\cdot 2}{x} = \frac{2.4}{x}\]

OpenStudy (anonymous):

Thanks, have a good one, PQ